Gel is two years older than Mel.In five years,Mel's age will be 34 years less than twice Mel's age will be.How old is each at present?
Answer by ankor@dixie-net.com(22740)

;et g = G's present age
let m = M's " "
:
Write an equation for each statement
:
Gel is two years older than Mel.
g = m + 2
or
m = g - 2
:
In five years, Mel's age will be 34 years less than twice Mel's age will be.
m + 5 = 2(g+5) - 34
m + 5 = 2g + 10 - 34
m = 2g - 24 - 5
m = 2g - 29
Replace m with (g-2)
g - 2 = 2g - 29
-2 + 29 = 2g - g
27 = g
G is 27 yrs old,
you can find M's age, check your solutions in the 2nd statement